Suppose U and V are subspace of R^n with U intersection V={0}. if {u1,.....uk} is a basis for U and {v1,.....vL} is a basis for V, prove that {u1.....uk,v1.......vL} is a basis for u+v.

1
Expert's answer
2021-12-28T12:44:14-0500

Let us prove that every vector in "U+V" can be written as a linear combination of  "\{u_1,…,u_k,v_1,…,v_l\}".


"U+V=\{u+v:\ u\in U,\ v\in V\}"

Since "\{u_1,…,u_k\}" is a basis of "U", we can write "u=\alpha_1 u_1+…+\alpha_ku_k" .

Since "\{v_1,…,v_l\}" is a basis of "V", we can write "v=\beta_1 v_1+…+\beta_lv_l" .

So, "u+v=\alpha_1u_1+…+\alpha _ku_k+\beta _1 v_1+…+\beta _lv_l" .


Now we need to prove that "\{u_1,…,u_k,v_1,…,v_l\}" is linearly independent.

Suppose we have "\alpha_1u_1+…+\alpha _ku_k+\beta _1 v_1+…+\beta _lv_l=0" .

It can be written as "\sum_{i=1}^k\alpha _iu_i=-\sum_{j=1}^l\beta_jv_j" .


We can see the left hand side is in "U" and the right hand side is in "V" . It means that both are in "U\cap V". Since "U\cap V=\{0\}" , it follows that "\sum_{i=1}^k\alpha _iu_i=\sum_{j=1}^l\beta_jv_j=0" .


Using the fact that "\{u_1,…,u_k\}" is a basis of "U" and "\{v_1,…,v_l\}" is a basis of "V" , we get that "\alpha_1=…=\alpha_k=0" and "\beta_1=…=\beta_l=0" .

So, "\{u_1,…,u_k,v_1,…,v_l\}" is linearly independent.


Now we can conclude that "\{u_1,…,u_k,v_1,…,v_l\}" is a basis for "U+V" .